# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

### v3

@@ Abstract
Integrating structured graph data with rich textual information from nodes poses a significant challenge, particularly for heterophilic node classification. Current approaches often struggle with computational costs or effective fusion of disparate modalities. We propose Graph Masked Language Model (GMLM), a novel architecture efficiently combining Graph Neural Networks (GNNs) with Pre-trained Language Models (PLMs). GMLM introduces three key innovations: (i) a dynamic active node selection strategy for scalable PLM text processing; (ii) a GNN-specific contrastive pretraining stage using soft masking with a learnable graph [MASK] token for robust structural representations; and (iii) a dedicated fusion module integrating RGCN-based GNN embeddings with PLM (GTE-Small & DistilBERT) embeddings. Extensive experiments on heterophilic benchmarks (Cornell, Wisconsin, Texas) demonstrate GMLM’s superiority. Notably, GMLM(DistilBERT) achieves significant performance gains, improving accuracy by over 4.7% on Cornell and over 2.0% on Texas compared to the previous best-performing baselines. This work underscores the benefits of targeted PLM engagement and modality-specific pretraining for improved, efficient learning on text-rich graphs.
@@ 1 Introduction
Graph Neural Networks (GNNs) [1, 2] have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes.. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT [3], have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning, especially for text-rich graphs, remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
While architectures like Graph Attention Networks (GATs) [4] offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) [5] can handle typed edges, they may not inherently utilize deep semantic understanding from node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we adapt and combine the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs, while addressing computational scalability ?
Traditional node feature masking in GNNs [6] often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
In this paper, we present Graph Masked Language Model (GMLM), a novel framework that bridges GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings and computational efficiency. Our approach introduces three key innovations:
@@
    A dynamic active node selection strategy that stochastically identifies a subset of nodes during each training iteration. Only the texts associated with these active nodes are processed by the PLM, and the primary task loss is computed only on these nodes, significantly reducing computational overhead.
@@
    A GNN-specific contrastive pretraining stage that employs a soft masking mechanism. This mechanism interpolates original node features with a learnable graph-specific [MASK] token embedding, guided by a dynamically generated perturbation mask. This stage pre-trains the GNN (an RGCN in our implementation) to learn robust structural representations before end-to-end fine-tuning.
@@
    A dedicated fusion network designed to effectively combine the comprehensive structural embeddings derived from the GNN path with the selectively generated, deep semantic embeddings from the PLM path, leading to rich, multi-modal node representations.
This combination allows our model to learn both structural and semantic patterns effectively, leading to more robust and contextually aware node representations, particularly beneficial for challenging heterophilic node classification tasks.
@@
    We propose a novel architecture, GMLM, that efficiently integrates a GNN (specifically an RGCN) with a PLM by introducing a dynamic active node selection mechanism for scalable text processing and a GNN-focused contrastive pretraining phase.
@@
    We introduce a GNN contrastive pretraining method utilizing a soft feature masking technique with a learnable graph [MASK] token. This pretrains the GNN to capture robust structural information and initializes the [MASK] token embedding meaningfully (Section 5.1).
@@
    We design a fusion strategy and an end-to-end fine-tuning process that utilizes the pre-trained GNN and selectively engages the PLM, allowing for effective learning from both graph structure and node texts for improved node classification, especially on heterophilic graphs (Section 5.2).
@@ 2 Related Works
Graph Neural Networks (GNNs) [2] provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) [4] integrated self-attention [7]. This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing [8], [9]. These Graph Transformers, reviewed systematically [10], can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models [3] that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) [11] framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure [12]; injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules [13]; interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information [14]; and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities [15]. We provide  Table 6 in  Appendix E giving an overview of the difference in methods and implementation.
@@ 3.2.1 Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) [5] extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
@@ 3.3 Pre-trained Language Models (PLMs)
Pre-trained Language Models (PLMs), such as BERT [3] and its variants ( DistilBERT [16]), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture [17] and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $[w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM (specifically DistilBERT) to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
@@ 3.4 Node Feature Masking in GNNs
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN [18, 19]. Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
@@ 3.5 Heterophily in Graphs
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar [20]. Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
@@ 3.6 Contrastive Learning
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss [21]:
@@
Figure 1: Overview of the GMLM architecture and its two-phase training procedure. (Phase A) GNN Contrastive Pre-Training: Two augmented views of the input graph features are generated using random masks and a learnable GNN [MASK] token, then processed by a shared RGCN stack. An NT-Xnet loss encourages similar representations for positive pairs. (Phase B) End-to-End Fine-Tuning: The graph and associated texts are input. Active Node Selection directs text processing for the PLM branch and guides GNN input feature soft masking. Outputs from the GNN and PLM branches are projected, fused, and passed to a classification head. The GNN is initialized from Phase A.
@@ 4 Methodology
We introduce GMLM (Graph Masked Language Model), a novel architecture designed to effectively integrate graph structural information with rich textual semantics from nodes for robust node representation learning, particularly targeting heterophilic graphs. The GMLM architecture is characterized by two main processing pathways for graph and text data, a specialized GNN input perturbation mechanism, and a dedicated fusion module. An overview of the entire pipeline is depicted in Figure 1.
@@ 4.1 GNN Input Feature Perturbation with Soft Masking
To enhance the robustness of the GNN and to provide a meaningful training signal during both the contrastive pretraining phase (detailed in Section 5.1) and the end-to-end fine-tuning phase, we employ a soft masking strategy for the input node features. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{x}}$ be the matrix of initial node features. Given a binary perturbation mask $\mathbf{m}_{pert}\in\{0,1\}^{N}$, which indicates the nodes whose features are to be perturbed, the soft masking operation produces perturbed features $\mathbf{X}^{\prime}$ as follows:
@@
where $\mathbf{x}_{i}$ is the original feature vector for node $v_{i}$, $\mathbf{x}^{\prime}_{i}$ is the perturbed feature vector, $\mathbf{e}_{mask}^{gnn}\in\mathbb{R}^{d_{x}}$ is a learnable GNN-specific mask token embedding, and $\beta\in[0,1]$ is a hyperparameter controlling the interpolation weight. This GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ is initialized using Xavier uniform initialization and learned alongside other model parameters. The perturbation mask $\mathbf{m}_{pert}$ is generated dynamically during training (details in Sections 5.1 and 5.2).
4.2 Graph Neural Network (GNN) Branch
[…]
To prepare for fusion with PLM embeddings, these raw GNN embeddings are projected to match the PLM’s hidden dimensionality $d_{PLM}$ using a linear projection layer:
$\mathbf{h}_{G,i}^{proj}=\mathbf{W}_{G}^{proj}\mathbf{h}_{G,i}+\mathbf{b}_{G}^{proj}$ (5)
where $\mathbf{W}_{G}^{proj}\in\mathbb{R}^{d_{PLM}\times d_{G}}$ and $\mathbf{b}_{G}^{proj}\in\mathbb{R}^{d_{PLM}}$ are learnable parameters. The activation for these projected embeddings occurs as part of the subsequent fusion network.
4.3 Pre-trained Language Model (PLM) Branch with Active Node Selection

### v4

@@ Abstract
Integrating powerful but computationally expensive Pre-trained Language Models (PLMs) with Graph Neural Networks (GNNs) is a key challenge, especially on text-rich heterophilic graphs. We propose the Graph Masked Language Model (GMLM), a framework designed for the efficient and effective fusion of graph structure and text semantics. GMLM employs a two-stage process: first, a contrastive pre-training stage with a novel soft masking technique builds a robust multi-scale GNN; second, an end-to-end fine-tuning stage uses a dynamic active node selection strategy for scalability and a bi-directional cross-attention module for deep fusion. Experiments on five heterophilic benchmarks show GMLM achieves state-of-the-art results on four, significantly outperforming prior GNN and large LLM-based methods. For instance, it improves accuracy on the Texas dataset by over 8% and on Wisconsin by nearly 5%. Our work demonstrates that a sophisticated, deeply-integrated architecture can be more effective and efficient than larger, general-purpose models for text-rich graph representation learning.
GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification
Aarush Sinha Independent Researcher aarush.sinha@gmail.com
@@ 1 Introduction
Graph Neural Networks (GNNs) Scarselli et al. (2009); Zhou et al. (2021) have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT Devlin et al. (2019), have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
While architectures like Graph Attention Networks (GATs) Veličković et al. (2018) offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) can handle typed edges, they do not inherently perform deep semantic reasoning on node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we develop a deeply integrated framework that combines the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs, while addressing computational scalability?
Traditional node feature masking in GNNs Mishra et al. (2021) often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
In this paper, we present a novel framework that deeply integrates GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings and computational efficiency. Our approach introduces three key innovations:
@@
    A dynamic active node selection strategy that stochastically identifies a subset of nodes during each training iteration. By processing only the texts of these active nodes with the PLM, we drastically reduce the computational overhead and memory footprint, making the approach scalable.
@@
    A GNN-specific contrastive pre-training stage that pre-trains a deep, multi-scale GNN (an RGCN in our implementation) to learn robust structural representations. This stage employs a novel soft masking mechanism that interpolates original node features with a learnable, graph-specific [MASK] token embedding, providing a more stable and informative training signal than hard masking.
@@
    A bi-directional cross-attention fusion module that moves beyond simple concatenation. It allows the multi-scale graph embeddings and the deep semantic PLM embeddings to mutually inform and refine one another, creating a rich, contextually-aware joint representation before the final classification.
This combination of selective PLM processing, robust GNN pre-training, and sophisticated fusion allows our model to learn both structural and semantic patterns effectively, leading to more powerful and contextually aware node representations for challenging node classification tasks.
@@
    We propose a novel and scalable framework that deeply integrates a multi-scale GNN with a PLM. The framework’s efficiency is achieved through a dynamic active node selection mechanism for on-demand text processing, making it tractable for large graphs.
@@
    We introduce a GNN-specific contrastive pre-training stage that utilizes a soft feature masking technique with a learnable graph [MASK] token. This stage is designed to effectively train a deep, multi-scale GNN architecture, allowing it to learn robust and hierarchical structural representations before end-to-end fine-tuning (Section 5.1).
@@
    We design a sophisticated bi-directional cross-attention fusion module that moves beyond simple concatenation. This mechanism allows the graph and text representations to mutually inform and refine one another, creating a richer, more contextually-aware joint representation for improved node classification, especially on challenging heterophilic graphs (Section 5.2).
@@ 2 Related Works
Graph Neural Networks (GNNs) Zhou et al. (2021) provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) Veličković et al. (2018) integrated self-attention Vaswani et al. (2017). This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing Ying et al. (2021), Yun et al. (2020). These Graph Transformers, reviewed systematically Shehzad et al. (2024), can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models Devlin et al. (2019) that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) Li et al. (2023a) framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure Plenz and Frank (2024); injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules Yuan and Färber (2024); interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information Henderson et al. (2023); and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities Ioannidis et al. (2022).
@@ 3.2.1 Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
@@ 3.3 Pre-trained Language Models (PLMs)
Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2019) and its variants ( DistilBERT Sanh et al. (2020)), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture Vaswani et al. (2023) and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $[w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM (specifically DistilBERT) to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
@@ 3.4 Node Feature Masking in GNNs
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN Hu et al. (2020); You et al. (2020). Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
@@ 3.5 Heterophily in Graphs
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar Zhu et al. (2020a). Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
@@ 3.6 Contrastive Learning
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss Sohn (2016):
@@
Figure 1: Caption
@@ 4 Methodology
We introduce a novel architecture designed to effectively integrate deep graph structural information with rich textual semantics from nodes for robust node representation learning. The architecture is characterized by two main processing pathways for graph and text data, a multi-scale GNN with residual connections, a sophisticated bi-directional cross-attention fusion module, and specialized input perturbation mechanisms to enhance robustness and training stability.
@@ 4.1 GNN Input Feature Perturbation with Soft Masking
To enhance the robustness of the GNN and to provide a meaningful training signal, we employ a soft masking strategy for the input node features. This is applied during both a contrastive pre-training phase and the final end-to-end fine-tuning. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{x}}$ be the matrix of initial node features for $N$ nodes. Given a binary perturbation mask $\mathbf{m}_{pert}\in\{0,1\}^{N}$, the soft masking operation produces perturbed features $\mathbf{X}^{\prime}$ as follows:
@@
where $\mathbf{x}_{i}$ is the original feature vector for node $v_{i}$, $\mathbf{e}_{mask}^{gnn}\in\mathbb{R}^{d_{x}}$ is a learnable GNN-specific mask token embedding, and $\beta\in[0,1]$ is a hyperparameter controlling the interpolation weight. This GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ is initialized using Xavier uniform initialization and learned alongside other model parameters.
4.2 Multi-Scale Relational GNN Branch
The GNN branch is designed as a deep, multi-scale architecture to capture structural information across different neighborhood depths. It utilizes a stack of four Relational Graph Convolutional Network (RGCN) layers. The input to the first RGCN layer is the perturbed node feature matrix $\mathbf{X}^{\prime}$ from Equation 4.
Each of the four GNN blocks consists of an RGCNConv layer, followed by GraphNorm, a GELU activation function, and Dropout. To facilitate the training of this deep architecture and combine features from different network depths, we incorporate residual connections with linear projections. The hidden representation $\mathbf{H}^{(l)}$ at layer $l$ is computed as:
$\mathbf{Z}^{(l)}$ $=\text{GELU}(\text{GraphNorm}(\text{RGCN}^{(l)}(\mathbf{H}^{(l-1)},\mathcal{E})))$ (5)
$\mathbf{H}^{(l)}$ $=\text{Dropout}(\mathbf{Z}^{(l)})+\text{Proj}^{(l-1)}(\mathbf{H}^{(l-2)})$ (6)
where $\mathbf{H}^{(0)}=\mathbf{X}^{\prime}$, $\mathcal{E}$ represents the graph edges and their types, and $\text{Proj}^{(l-1)}$ is a linear projection. This structure allows the model to maintain and propagate information from earlier layers. We employ gradient checkpointing [^1] Specifically, torch.utils.checkpoint.checkpoint is used for each GNN block to trade a small amount of re-computation for significant memory savings, enabling the use of a deeper GNN. on each GNN block to manage memory consumption during backpropagation.
Instead of using only the output of the final GNN layer, we adopt a multi-scale fusion strategy to create a more comprehensive graph representation. The embeddings from all four GNN layers, $\{\mathbf{H}^{(1)},\mathbf{H}^{(2)},\mathbf{H}^{(3)},\mathbf{H}^{(4)}\}$, are collected. These are then combined using a learnable weighted sum after being projected to the PLM’s hidden dimension, $d_{PLM}$:
$\mathbf{h}_{G,i}=\text{LayerNorm}\left(\sum_{l=1}^{4}w_{l}\cdot\text{Proj}^{(l)}(\mathbf{h}_{i}^{(l)})\right)$ (7)
where $\mathbf{h}_{i}^{(l)}$ is the embedding for node $v_{i}$ from layer $l$, $\text{Proj}^{(l)}$ is a linear projection to dimension $d_{PLM}$, and $w_{l}$ are learnable weights derived from a softmax-normalized parameter vector. The final graph embedding for node $v_{i}$ is $\mathbf{h}_{G,i}\in\mathbb{R}^{d_{PLM}}$.
4.3 Pre-trained Language Model (PLM) Branch